### v1

# XY-Cut++: Advanced Layout Ordering via Hierarchical Mask Mechanism on a Novel Benchmark

Abstract
Document Reading Order Recovery is a fundamental task in document image understanding, playing a pivotal role in enhancing Retrieval-Augmented Generation (RAG) and serving as a critical preprocessing step for large language models (LLMs). Existing methods often struggle with complex layouts(e.g., multi-column newspapers), high-overhead interactions between cross-modal elements (visual regions and textual semantics), and a lack of robust evaluation benchmarks. We introduce XY-Cut++, an advanced layout ordering method that integrates pre-mask processing, multi-granularity segmentation, and cross-modal matching to address these challenges. Our method significantly enhances layout ordering accuracy compared to traditional XY-Cut techniques. Specifically, XY-Cut++ achieves state-of-the-art performance (98.8 BLEU overall) while maintaining simplicity and efficiency. It outperforms existing baselines by up to 24% and demonstrates consistent accuracy across simple and complex layouts on the newly introduced DocBench-100 dataset. This advancement establishes a reliable foundation for document structure recovery, setting a new standard for layout ordering tasks and facilitating more effective RAG and LLM preprocessing.
1 Introduction
Document Reading Order Recovery is a fundamental task in document image understanding, serving as a cornerstone for enhancing Retrieval-Augmented Generation (RAG) systems and enabling high-quality data preprocessing for large language models (LLMs). Accurate layout recovery is essential for applications such as digital publishing and knowledge base construction. However, this task faces several significant challenges: (1) complex layout structure (e.g., multi-column layout, nested text boxes, non-rectangular text regions, cross-page content), (2) inefficient cross-modal alignment due to the high computational costs of integrating visual and textual features, and (3) the lack of standardized evaluation protocols for block-level reading order. Traditional approaches like XY-Cut [2] fail to model semantic dependencies in complex designs, while deep learning methods such as LayoutReader [13] suffer from prohibitive latency, limiting real-world deployment. Compounding these issues, existing datasets like ReadingBank [13] focus on word-level sequence annotations, which inadequately evaluate block-level structural reasoning—a necessity for real-world layout recovery. Although OmniDocBench [8] recently introduced block-level analysis support, its coverage of diverse complex layouts (e.g., newspapers, technical reports) remains sparse, further hindering systematic progress.
To address these challenges, we propose XY-Cut++, an advanced framework for layout ordering that incorporates three core innovations: (a) pre-mask processing to mitigate interference from high-dynamic-range elements (e.g., title, figure, table), (b) multi-granularity region splitting for the adaptive decomposition of diverse layouts, and (c) lightweight cross-modal matching leveraging minimal semantic cues. Our approach significantly outperforms traditional methods in layout ordering accuracy. Additionally, we introduce DocBench-100, a novel benchmark dataset designed for evaluating layout ordering techniques. It includes 100 pages (30 complex and 70 regular layouts) with block-level reading order annotations. Extensive experiments on DocBench-100 demonstrate that our method achieves state-of-the-art performance. Specifically, our method achieves BLEU-4 scores of 98.6(complex) and 98.9(regular), surpassing baselines by 24% on average while maintaining 1.06× FPS of geometric-only approaches.
The contributions of this paper are summarized as follows:
  1. 1.
    We present a simple yet high-performance enhanced XY-Cut framework that fuses shallow semantics and geometry awareness to achieve accurate layout ordering.
  2. 2.
    We establish a new benchmark dataset for layout ordering by curating complex layouts from document detection datasets, enriching the diversity of challenging layouts and offering a more comprehensive resource than existing datasets.
  3. 3.
    We achieved state-of-the-art results on both existing and new datasets, and provided extensive ablations to each component in our methods.
Figure 2: XY-Cut Recursive Partitioning Workflow and Failure Analysis in Complex Layouts: (1) Initial document segmentation steps, (2) Connectivity assumption violations caused by cross-layout cell structures (cell 5), and (3) Error propagation through subsequent layout ordering.
2 Related Work
Document Reading Order Recovery has seen significant advancements, evolving from early heuristic rule-based approaches to sophisticated deep learning approaches.
2.1 Traditional Approaches
The XY-Cut algorithm [2] is a foundational technique in Document Reading Order Recovery that recursively divides documents into smaller regions based on horizontal and vertical projections. As illustrated in Fig. 2, while it is effective for simple layouts, it struggles with complex structures, leading to inaccuracies in reading order recovery. Specifically, the rigid threshold mechanisms of XY-Cut can introduce hierarchical errors when handling intricate layouts, resulting in suboptimal performance. To address these limitations, various enhancements have been proposed. For instance, dynamic programming optimizations [7] have been introduced to improve segmentation stability. Additionally, mask-based normalization techniques [11] have been developed to mitigate some of the challenges associated with complex layouts. However, these improvements are still insufficient for handling intricate layouts and cross-page content. In our analysis of the XY-Cut algorithm, as discussed in [7, 11] and illustrated in Fig. 3, we identified that many challenging cases arise from L-shaped inputs. A straightforward solution involves initially masking these L-shaped regions and subsequently restoring them. Upon implementing this approach, we found it to be remarkably effective while maintaining both simplicity and efficiency.
Figure 3: Challenges posed by L-shaped inputs: (1) segmentation failure due to the inability to process L-shaped structures, and (2) missegmentation caused by improper handling of L-shaped regions. ( The correct order is ➁+➂ ➀)
2.2 Deep Learning Approaches
Recent advances in Document Reading Order Recovery have been driven by deep learning models that effectively leverage multimodal cues. The LayoutLM series [14, 17, 3, 16, 1] pioneered the integration of textual semantics with 2D positional encoding, where LayoutLMv2 [17] introduced spatially aware pre-training objectives for cross-modal alignment and LayoutLMv3 [3] further unified text/image embeddings through self-supervised masked language modeling and image-text matching. LayoutXLM [16] extended this framework to multilingual document understanding. XYLayoutLM [1] advanced the field with its Augmented XY-Cut algorithm and Dilated Conditional Position Encoding, addressing variable-length layout modeling and generating reading order-aware representations for enhanced document understanding. Building upon these foundations, LayoutReader [13] demonstrated the effectiveness of deep learning in explicit reading order prediction through the sequential modeling of paragraph relationships. Critical to these advancements are large-scale datasets like DocBank [5], which provides weakly supervised layout annotations for fine-grained spatial modeling, and PubLayNet [20], which contains over 360,000 scientific document pages with hierarchical labels that encode implicit reading order priors through structural patterns. Specialized benchmarks like TableBank [4] further address domain-specific challenges by preserving the ordering of tabular data for table structure recognition.
Architectural innovations have significantly enhanced spatial reasoning capabilities. DocRes [19] introduces dynamic task-specific prompts (DTSPrompt), enabling the model to perform various document restoration tasks, such as dewarping, deshadowing, and appearance enhancement, meanwhile improving the overall readability and structure of documents.  [18] iteratively aggregates features across different layers and resolutions, resulting in more refined feature representations that are beneficial for complex layout analysis tasks. By leveraging Hierarchical Document Analysis (HDA), models can more effectively capture intricate structural relationships within documents, facilitating more accurate predictions of reading order. Furthermore, advancements in unsupervised learning methods, such as those employed in DocUNet [6], have enabled more effective handling of document distortions, thereby enhancing OCR performance and layout analysis accuracy. DocFormer [15] unified text, visual, and layout embeddings through transformer fusion, improving contextual understanding for logical flow prediction. Complementary to these approaches, EAST [21] established robust text detection through geometry-aware frameworks, serving as a critical preprocessing step for element-level sequence derivation.
2.3 Benchmarks for Document Reading Order Recovery
Document Reading Order Recovery has seen significant advancements with deep learning models like LayoutLM [14] and LayoutReader [13], which integrate visual and textual features for tasks such as reading order prediction. However, existing methods face critical limitations in handling complex layouts (e.g., multi-column structures, cross-page content). A key challenge is the lack of benchmarks that directly evaluate block-level reading order, which is essential for applications like document digitization. While datasets such as ReadingBank [13] provide word-level annotations for predicting reading sequences, their design complicates the evaluation of methods focused on block-level performance. Specifically, word-level sequence annotations cannot simplify the assessment of dependencies between text blocks (e.g., the order of paragraphs or figures spanning multiple columns), which are essential for modeling complex layouts. Recently, OmniDocBench [8] has supported block-level analysis but suffers from limited coverage of complex layouts and sparse representation of domain-specific layouts (e.g., newspapers). To address these gaps, we introduce DocBench-100, which offers a broader range of layout structures and explicit metrics for assessing reading order accuracy, thereby enabling robust benchmarking of layout recovery systems.
3 Methodology
Our geometry-semantic fusion pipeline, as shown in Fig. 4, consists of four coordinated stages: (1) PP-DocLayout [10] extracts visual features and shallow semantic labels from document images; (2) highly dynamic elements (e.g., titles, tables, figures) are pre-masked to alleviate the ”L-shape” problem; (3) cross-layout elements are identified through global layout analysis, followed by mask processing, real-time density estimation, and heuristic sorting to ensure logical content ordering; and (4) masked elements are re-mapped using nearest IoU edge-weighted margins. By integrating geometry-aware processing with shallow semantic information, our method achieves state-of-the-art performance on DocBench-100 and OmniDocBench [8], demonstrating its effectiveness in addressing complex document layout challenges.
Figure 4: End-to-End Layout Ordering for Diverse Document Layouts Framework Overview: (a) Layout Detection(We use PP-Doclayout[10]), (b) Pre-Mask Processing, (c) Multi-Granularity Segmentation, and (d) Cross-Model Matching.
Figure 5: Multi-Granularity Segmentation: (1) cross-layout element masking, (2) preliminary segmentation via pre-cut, and (3) ecursive density-driven partitioning. The enhanced XY-Cut algorithm adapts its splitting axis selection through real-time density evaluation ($\tau_{d}$), prioritizing horizontal splits for content-dense regions and vertical splits otherwise.
Figure 6: Cross-Modal Matching: (1) semantic hierarchy-aware stage decomposition and (2) adaptive distance metric matching with dynamic policies and semantic-specific tuning. Subsequently, cells are reordered based on Index, Label Priority, Y1, and X1.
3.1 Multi-Granularity Segmentation
As shown in Fig. 5, our hybrid segmentation framework combines masking, preliminary splitting, and layout-aware refinement through three key phases:
Phase 1: Cross-Layout Detection Compute document-level median bounding box length and establish adaptive threshold with scaling factor $\beta=1.3$:
$\mathcal{T}_{l}=\beta\cdot\mathrm{median}(\{l_{i}\}_{i=1}^{N})$ (1)
Detect cross-layout elements using dual criteria:
$\mathcal{C}_{\mathrm{cross}}(B_{i})=\begin{cases}T&l_{i}>\mathcal{T}_{l}\ \land\ \sum_{j\neq i}\mathbb{I}_{\mathrm{overlap}}(B_{i},B_{j})\geq 2\\ F&\text{otherwise}\end{cases}$ (2)
Masked cross-layout elements are preserved for subsequent layout-aware splitting, while standard components undergo immediate processing.
Phase 2: Geometric Pre-Segmentation Central elements and isolated graphical components are identified through:
$\mathcal{P}(B_{i})=\mathbb{I}\left(\frac{\|c_{i}-c_{\text{page}}\|_{2}}{d_{\text{page}}}\leq 0.2\right)\land(\phi_{\text{text}}(B_{i})=\infty)$ (3)
where $c_{i}$ denotes bounding box center coordinates, and classified visual elements include {Figure, Table}. Here,$\phi_{\text{text}}(B_{i})$ represents the distance from the bounding box $B_{i}$ to the nearest text box,and $\infty$ indicates that $B_{i}$ is not adjacent to any text box.
Phase 3: Density-Driven Refinement The adaptive XY-Cut algorithm dynamically selects splitting axes based on regional content density:
$\tau_{d}=\frac{\sum_{k=1}^{K_{c}}w_{k}^{(C_{c})}h_{k}^{(C_{c})}}{\sum_{k=1}^{K_{s}}w_{k}^{(C_{s})}h_{k}^{(C_{s})}}$ (4)
Direction-aware splitting strategy:
$\mathcal{S}(R)=\begin{cases}\mathrm{XY\text{-}Cut}&\tau_{d}>\theta_{v}\ (\theta_{v}=0.9)\\ \mathrm{YX\text{-}Cut}&\text{otherwise}\end{cases}$ (5)
Atomic region representation:
$R_{i}=\langle x_{1}^{(i)},y_{1}^{(i)},x_{2}^{(i)},y_{2}^{(i)},C_{i}\rangle,\quad C_{i}\in\mathcal{C}_{\mathrm{type}}$ (6)
with $(x_{1}^{(i)},y_{1}^{(i)})$ and $(x_{2}^{(i)},y_{2}^{(i)})$ defining the bounding box coordinates. The content type $C_{i}$ distinguishes between Cross-layout (spanning multiple grid units) and Single-layout (contained within one grid unit) components.
3.2 Cross-Modal Matching
To establish cross-modal reading coherence, we propose a geometry-aware alignment framework with two core components in Fig. 6.
Multi-Stage Semantic Filtering: Orders candidates through label priority:
$\begin{gathered}\mathcal{L}_{\text{order}}:\mathcal{L}_{\text{cross-layout}}\succ\mathcal{L}_{\text{title}}\succ\mathcal{L}_{\text{vision}}\succ\mathcal{L}_{\text{others}}\\ \mathcal{M}_{\text{sem}}^{(l)}=\left\{(B_{\text{p}},B_{\text{o}}^{\prime})\Bigg{|}\begin{aligned} &L_{\text{p}}=l\\ &\land\ L_{\text{o}}^{\prime}\succeq l\end{aligned}\right\},\quad l\in\mathcal{L}_{\text{order}}\\ \mathcal{M}_{\text{sem}}=\bigcup_{l\in\mathcal{L}_{\text{order}}}\mathcal{M}_{\text{sem}}^{(l)}\end{gathered}$ (7)
where $\mathcal{L}_{\text{vision}}$ denotes visual element detection objectives including tables, images, seals, and their corresponding titles, with $B_{p}$ representing pending bounding boxes and $B^{\prime}_{o}$ indicating ordered candidate regions.
Adaptive Distance Metric: Computes joint geometric distance with early termination. Given layout element $B_{\text{p}}=(x_{1},y_{1},x_{2},y_{2})$ and candidate $B_{\text{o}}^{\prime}=(x_{1}^{\prime},y_{1}^{\prime},x_{2}^{\prime},y_{2}^{\prime})$, we define:
$D(B_{\text{p}},B_{\text{o}}^{\prime},l)=\sum_{k=1}^{4}w_{k}\cdot\phi_{k}(B_{\text{p}},B_{\text{o}}^{\prime})$ (8)
where $\phi_{k}$ encodes four geometric constraints:
  • Intersection Constraints:
    $\phi_{1}=\begin{cases}1,&\text{if }\mathrm{dir}(B_{p})\neq\mathrm{dir}(B_{o}^{\prime})\lor\mathrm{IoU}_{\text{partial}}<\tau_{\text{overlap}}\\ 0,&\text{otherwise}\end{cases}$
  • Boundary Proximity:
    $\phi_{2}=w_{\text{edge}}\times\begin{cases}d_{x}+d_{y},&\text{(diagonal adjacency)}\\ \min(d_{x},d_{y}),&\text{(axis-aligned)}\end{cases}$
  • Vertical Continuity:
    $\phi_{3}=\begin{cases}-y_{2}^{\prime},&l\in\mathcal{L}_{\text{cross-layout}}\land y_{1}>y_{2}^{\prime}\\ y_{1}^{\prime},&\text{(baseline alignment)}\end{cases}$
  • Horizontal Ordering:
    $\phi_{4}=x_{1}^{\prime}$
We further introduce two parameterization strategies:
Dynamic Weight Adaptation. Scale-sensitive weights via dimensional analysis:
$\boldsymbol{w}_{k}=[\max(h,w)^{2},\max(h,w),1,\max(h,w)^{-1}]$ (9)
where h, w denote page dimensions, it establishes a distance metric with clear priorities.
Figure 7: Visualization of Complex Page $D_{c}$ from DocBench-100 Dataset Using Different Layout Analysis Methods. (a) Input Image, (b) XY-Cut[2], classic projection-based segmentation, (c) MinerU[12], an end-to-end Document Content Extraction tool, (d) XY-Cut++, our proposed method.
Semantic-Specific Tuning. Optimal weights from grid search on 2.8k documents:
$\boldsymbol{w}_{\text{edge}}=\begin{cases}[1,0.1,0.1,1]&l\in\mathcal{L}_{\text{title}}\cap\mathcal{O}_{\text{horiz}}\\ [0.2,0.1,1,1]&l\in\mathcal{L}_{\text{title}}\cap\mathcal{O}_{\text{vert}}\\ [1,1,0.1,1]&l\in\mathcal{L}_{\text{cross-layout}}\\ [1,1,1,0.1]&l\in\mathcal{L}_{\text{otherwise}}\end{cases}$ (10)
Statistical validation demonstrates significant improvements: +2.3 BLEU-4 over uniform baselines. The overall process can be simply expressed in algorithm 1.
Algorithm 1 The overall flow of cross-model matching
1:Initialize $\mathcal{L}_{\text{order}}$: $\mathcal{L}_{\text{cross-layout}}\succ\mathcal{L}_{\text{title}}\succ\mathcal{L}_{\text{vision}}\succ\mathcal{L}_{\text{others}}$
2:Initialize Label Priority: doc title ¿ title ¿ otherwise
3:Initialize: $\mathcal{M}_{\text{aligned}}\leftarrow\mathcal{B}_{\text{ordered}}^{\prime}$
4:for each masked label $L\in\mathcal{L}_{\text{order}}$ do
5:     for each pending box $B\in\mathcal{B}_{\text{pending-L}}$ do
6:         Initialize $D_{\min}\leftarrow\infty$
7:         for each ordered box $B^{\prime}\in\mathcal{B}_{\text{ordered}}^{\prime}$ do
8:              $D_{\text{curr}}\leftarrow 0$
9:              for $k\leftarrow 1$ to 4 do
10:                  $D_{\text{curr}}\leftarrow D_{\text{curr}}+w_{k}\cdot\phi_{k}(B,B^{\prime})$
11:                  if $D_{\text{curr}}>D_{\min}$ then
12:                       break $\triangleright$ filter out
13:                  end if
14:              end for
15:              if $D_{\text{curr}}<D_{\min}$ then
16:                  $D_{\min}\leftarrow D_{\text{curr}}$, $B^{\prime}_{\text{best}}\leftarrow B^{\prime}$
17:              end if
18:         end for
19:         $\mathcal{M}_{\text{aligned}}\leftarrow\mathcal{M}_{\text{aligned}}\cup\{(B,B^{\prime}_{\text{best}})\}$
20:     end for
21:     Reorder $\mathcal{M}_{\text{aligned}}$ by (Index, Label Priority, Y1, X1)
22:end for
4 Experiments
Figure 8: Visualization of Regular Page $D_{r}$ from DocBench-100 Dataset Using Different Layout Analysis Methods. (a) Input Image, (b) XY-Cut[2], a classic projection-based segmentation, (c) MinerU[12], an end-to-end Document Content Extraction tool, (d) XY-Cut++, our proposed method.
Table 1: Progressive Component Analysis on DocBench-100.Metric Key: BLEU-4 $\uparrow$ / ARD $\downarrow$ / Tau $\uparrow$.
Method $D_{c}$ $D_{r}$ $\mu$
BLEU-4 $\uparrow$ ARD $\downarrow$ Tau $\uparrow$ BLEU-4 $\uparrow$ ARD $\downarrow$ Tau $\uparrow$ BLEU-4 $\uparrow$ ARD $\downarrow$ Tau $\uparrow$
XY-Cut 0.749 0.233 0.878 0.819 0.098 0.912 0.797 0.139 0.902
+Pre-Mask 0.818 0.196 0.887 0.823 0.087 0.920 0.822 0.120 0.910
+MGS 0.946 0.164 0.969 0.969 0.036 0.985 0.962 0.074 0.980
+CMM 0.986 0.023 0.995 0.989 0.003 0.997 0.988 0.009 0.996
Table 2: Ablation Study of Pre-Mask, Multi-Granularity Segmentation (MGS), and Cross-Modal Matching (CMM) on DocBench-100.Metric Key: BLEU-4 $\uparrow$.
Method Mask Mask Cross-Layout Pre-Cut Adaptive Scheme $\phi_{1}$ $\phi_{2}$ $\phi_{3}$ $\phi_{4}$ Dynamic Weights Multi-Stage BLEU-4 $\uparrow$ Baseline 0.797 +Pre-Mask 0.822 +MGS 0.905 0.914 0.923 0.962 +CMM 0.963 0.765 0.858 0.985 0.881 0.694 0.988
Table 3: Reading Order Recovery Performance: BLEU-4 $\uparrow$ Results on DocBench-100.The best results are in bold.
Method $D_{c}$ $D_{r}$ $\mu$
XY-Cut [2] 0.749 0.818 0.797
LayoutReader [3, 13] 0.656 0.844 0.788
MinerU [12] 0.701 0.946 0.873
XY-Cut++(ours) 0.986 0.989 0.988
Table 4: Reading Order Recovery Performance on Textual Content of OmniDocBench (Excluding Figures/Tables with Insignificant Layout Impact). Metric Key: BLEU-4 $\uparrow$ / ARD $\downarrow$ / Tau $\uparrow$. The red color indicates the best results and the blue color indicates the second-best results.
Method Single Double Three Complex Mean BLEU-4 ARD Tau BLEU-4 ARD Tau BLEU-4 ARD Tau BLEU-4 ARD Tau BLEU-4 ARD Tau XY-Cut [2] 0.895 0.042 0.931 0.695 0.230 0.794 0.702 0.090 0.923 0.717 0.120 0.866 0.753 0.118 0.878 LayoutReader [13, 3] 0.988 0.004 0.995 0.831 0.084 0.918 0.595 0.208 0.805 0.716 0.116 0.864 0.783 0.099 0.906 MinerU [12] 0.961 0.025 0.969 0.933 0.037 0.971 0.923 0.042 0.965 0.887 0.050 0.932 0.926 0.039 0.959 XY-Cut++(ours) 0.993 0.004 0.996 0.951 0.027 0.974 0.967 0.033 0.984 0.901 0.064 0.942 0.953 0.037 0.972
Table 5: Model Efficiency and Semantic Information Usage on DocBench-100 and OmniDocBench. Key Metrics: FPS (Total Pages/Total Times) $\uparrow$. FPS values are averaged over 10 runs on Intel(R) Xeon(R) Gold 6326 CPU @ 2.90GHz with 256GB memory. The red color indicates the best results and the blue color indicates the second-best results.
Method Semantic Info FPS DocBench-100 OmniDocBench Mean XY-Cut [2] 685 289 487 LayoutReader [3, 13] 17 27 22 MinerU [12] 10 12 11 XY-Cut++(ours) 781 248 514
In this section, we describe the experimental setup for evaluating our proposed method on the DocBench-100 benchmark. We compare our approach with several state-of-the-art baselines and demonstrate its effectiveness through both quantitative and qualitative analyses.
4.1 Dataset
We introduce DocBench-100, a novel benchmark dataset constructed from the document detection dataset [10], to address the lack of a standardized dataset for evaluating Document Reading Order Recovery tasks. Existing benchmarks fail to adequately cover the diverse layout complexities encountered in real-world documents. DocBench-100 consists of two subsets: the complex subset ($D_{c}$) and the regular subset ($D_{r}$), reflecting varying levels of layout complexity found in everyday documents.
Complex Subset ($D_{c}$): This subset contains 30 pages sourced from newspapers, books, and textbooks. The column distribution is 3.3% single-column, 6.7% double-column, 90% multi-column layouts (with $\geq 3$ columns and irregular document titles). These features make $D_{c}$ particularly challenging and representative of complex, multi-column layouts typical of newspapers and scientific journals.
Regular Subset ($D_{r}$): This subset contains 70 pages sourced from academic papers, books, textbooks, exam papers, slides, financial reports, notebooks, and magazines. The column distribution is 38.6% single-column, 54.4% double-column, and 7% three-column layouts. This subset represents more common and simpler layouts typically found in academic papers and books.
Dataset Composition: To reflect the prevalence of simpler layouts in daily use, we set the ratio of $D_{r}$ to $D_{c}$ at 7:3. The entire dataset comprises 100 original page images, each accompanied by a ground truth (GT) JSON file and an input JSON (No Index) file. The GT files contain essential information, including:
  • page_id: A unique identifier for each page.
  • page_size: The dimensions of the page.
  • bbox: Bounding boxes for each element on the page.
  • label: Class labels for different types of content (e.g., text, figure).
  • index: The reading order index for each element.
All reading orders have been manually verified to ensure their accuracy.
Necessity of DocBench-100: The introduction of DocBench-100 fills a critical gap in the field by providing a comprehensive dataset that covers both simple and complex layouts. This diversity allows for robust evaluation of reading order recovery models across various scenarios. The inclusion of detailed ground truth (GT) annotations facilitates direct testing and comparison with other models, thereby promoting reproducibility and advancing further research in this domain.
4.2 Setup
Baselines
We compare our method with the following state-of-the-art approaches:
  • XY-Cut [2]: A classic method for projection-based segmentation.
  • LayoutReader [3, 13]: A LayoutLMv3-based model fine-tuned on 500k samples.
  • MinerU [12]: An end-to-end document content extraction tool.
Evaluation Metrics
We evaluate the performance using the following metrics:
  • BLEU-4 [9] ($\uparrow$): Measures the similarity between candidate and reference texts using up to 4-gram overlap.
  • ARD ($\downarrow$): Absolute Relative Difference, quantifies prediction accuracy by comparing predicted and actual values.
  • Tau ($\uparrow$): Kendall’s Tau, measures the rank correlation between two sets of data.
  • FPS ($\uparrow$): Frames Per Second, a measure of how many frames a system can process per second.
4.3 Main Results
We evaluate our method on DocBench-100, analyzing component contributions and comparing against state-of-the-art baselines. All metrics are computed over the union of $D_{c}$ and $D_{r}$, subsets unless specified.
4.3.1 Progressive Component Ablation
Table 1 demonstrates the cumulative impact of each technical component:
  • XY-Cut Baseline: Achieves 0.749 BLEU-4 on complex layouts ($D_{c}$), showing limitations in handling complex-layout(e. g. L-shape) elements.
  • +Mask: Improves BLEU-4 by 6.9 points (from 0.749 to 0.818) on $D_{c}$ via adaptive thresholding (Equation (1), $\beta=1.3$), reducing false splits by 15.9%.
  • +MGS: Delivers a 19.7 absolute BLEU-4 gain on $D_{c}$ through three-phase segmentation, with density-aware splitting (Equation (5)) reducing ARD by 29.7%.
  • +CMM: Achieves near-perfect alignment (0.995 $\tau$) on $D_{c}$ through geometric constraints (Equations (7)-(10)), finalizing a 90.1% ARD reduction from baseline.
The complete model reduces ARD by 93.5% compared to baseline (0.139→0.009), demonstrating superior ranking consistency. Notably, our method maintains balanced performance across both subsets (0.988 $\mu$-BLEU), proving effective for diverse layout types.
4.3.2 Architectural Analysis
We perform a systematic dissection of core components through controlled ablations to evaluate their contributions:
Pre-Mask Processing (Pre-Mask): To alleviate the ”L-shaped” problem, we applied preliminary masking on highly dynamic elements such as titles, tables, and figures. This approach reduced visual noise and improved reading order recovery, resulting in a 2.5-point BLEU-4 score increase, as shown in Table 2.
Multi-Granularity Segmentation (MGS): Mask cross-layout is a very direct method to solve the problem that XY-Cut cannot segment L-shaped input. Pre-Cut realizes preliminary sub-page division through page analysis, thereby avoiding page content mixing affecting sorting. The adaptive splitting strategy enables reasonable segmentation through real-time density estimation. Table 2 shows additive benefits of the mask cross-layout (+8.3), Pre-Cut (+9.2), and the adaptive splitting scheme (+10.1) over the baseline (82.2 BLEU-4).
Cross-Modal Matching (CMM): As shown in Table 2, the single-stage strategy performs comparably to the baseline approach. This suggests that employing a detection model equipped with text, title, and annotation labels is sufficient to achieve nearly consistent performance across various scenarios. Notably, on the OmniDocBench dataset, which has few label categories, our method still achieves state-of-the-art results. Furthermore, we observe that the edge-weighted margin distance plays a crucial role among the four distance metrics examined. This finding highlights the significance of dynamic weights based on shallow semantics. In contrast, implementing cross-modal matching results in an overall improvement of 2.7 points in BLEU-4 scores.
4.3.3 Benchmark Comparison
As shown in Table 3, our approach establishes new benchmarks across all evaluation dimensions. Specifically, it outperforms XY-Cut by a significant margin, achieving a +23.7 absolute improvement on $D_{c}$ (from 74.9 to 98.6). Additionally, it surpasses LayoutReader by +5.3 on $D_{r}$ (from 94.6 to 98.9), despite not using any learning-based components. Furthermore, our method achieves a Kendall’s $\tau$ of 0.996 overall, indicating near-perfect ordinal consistency (with $p<0.001$ in the Wilcoxon signed-rank test). Visual results presented in Fig. 7 and 8 further demonstrate the robustness in handling multi-column layouts and cross-page elements, where previous methods frequently fail.
To further validate the versatility and robustness of our proposed method, we conducted extensive evaluations on the OmniDocBench dataset [8], which features a diverse and challenging set of document images. As shown in Table 4, our proposed method (XY-Cut++) achieves state-of-the-art (SOTA) performance across almost all layout types, despite challenging subpage nesting patterns (see Limitations). Notably, as shown in Table 5, XY-Cut++ achieves a superior balance between performance and speed, attaining an average FPS of 514 across DocBench-100 and OmniDocBench. This performance surpasses even the direct projection-based XY-Cut algorithm, which achieves an average FPS of 487. The significant speed improvement of XY-Cut++ is primarily attributed to semantic filtering, which minimizes redundant processing by handling each block only once. In contrast, XY-Cut requires repeatedly partitioning blocks into different subsets, resulting in increased recursive depth and computational overhead. This optimization enhances computational efficiency without losing performance, making XY-Cut++ more robust and versatile for diverse document layouts.
5 Conclusion
In this work, we address the critical challenges in document layout recovery and semantic conversion, particularly focusing on the needs of Retrieval-Augmented Generation (RAG) and Large Language Models (LLMs) for accurate information extraction. We introduce a hierarchical document parsing framework that incorporates a Mask mechanism-based layout recovery algorithm. This framework enables precise reading order recovery through pre-mask processing, multi-granularity segmentation, and cross-modal matching. Additionally, it achieves superior performance on the DocBench-100 and OmniDocbench datasets, outperforming baseline methods and providing robust support for downstream RAG and LLM tasks. It is noteworthy that while traditional methods are known for their computational efficiency, they often fall short in performance. Our approach not only maintains simplicity and efficiency but also matches or surpasses the performance of deep learning methods. This advantage is primarily attributed to two key innovations: (1) the strategic use of shallow semantic labels as structural priors during layout analysis, and (2) a hierarchical mask mechanism that effectively captures the document’s topological structure. These innovations not only enhance performance but also maintain model efficiency, suggesting promising directions for improving deep learning models in layout recovery tasks. We believe our work can inspire new approaches to tackle the challenging tasks in document understanding. Our code will be released at:https://github.com/liushuai35/PaddleXrc.
6 Limitations and Future Work
Our method currently struggles to handle pages with nested semantic structures, like sub-pages within a single page, as it cannot effectively sort elements within these complex layouts. To tackle this, we plan to enhance our framework by adding a sub-page detection module. This module will identify sub-page regions, enabling element sorting within them while maintaining the natural reading order (top-to-bottom, left-to-right) for the sub-pages themselves. By balancing global and local sorting, this approach aims to enhance the robustness of our method for complex page structures.
Figure 1: XY-Cut++ Architecture with Hierarchical Mask Mechanism and DocBench-100 Benchmark. (left) DocBench-100 dataset composition (9 document categories, Multi-type layout) and dual evaluation protocols: end-to-end image-based assessment (requiring layout parsing models) and direct JSON-based metric computation. (right) Algorithm workflow integrating adaptive pre-mask processing, multi-granularity segmentation, and cross-modal matching. Unlike neural network models, our method is simple, effective, and strongly interpretable.